OpenStudy (k.binks):

Find the distance between points (2, 4) and (5, 7)?

OpenStudy (joannablackwelder):

Can you draw the two points on a x-y graph?

OpenStudy (princeharryyy):

sqrt((5-2)^2+(7-4)^2) = sqrt ((3)^2+(3)^2) = 3sqrt(2) where sqrt = square root

OpenStudy (mrnood):

 You can see that you can make a right angle triangle from the 2 points. The length of the side shown as a is the difference in the y values of the points (7-4) The length of the side shown as b is the difference in the y values of the points (5-2) Then you can use Pythagoras to calculate the length of th ethe side c (the hypotenuse) a^2 + b^2 = c^2 this is the basis of the answer that @princeharryyy gave above